OpenStudy (anonymous):

I have this trig. illustration but I dont know how to solve and what formula

OpenStudy (imstuck):

Trig is awesome! Whatcha got?

OpenStudy (anonymous):



OpenStudy (anonymous):

I don't know what formula what to use and how to solve :(

OpenStudy (imstuck):

Ok, let's start with this... do you know what supplementary angles are?

OpenStudy (anonymous):

Yes, they add up to 180 deg right?? I know they must be equal to 180 deg but two given are exterior angle and one is interior

OpenStudy (anonymous):

so youre saying that (4x-3=180)?

OpenStudy (imstuck):

Is this an equilateral triangle, first off

OpenStudy (anonymous):

I dont think so

OpenStudy (imstuck):

Ok, I'm working something now...give me a sec

OpenStudy (anonymous):

Okay thank you very much!

OpenStudy (imstuck):

Ok here's what I came up with. Angle A is supplemental with 4x - 3 (that's what I am naming it, at least); angle B is supplemental with 9x + 12; and the other angle is 6x + 3. In order to find out what angle A is we need to subtract 4x - 3 from 180. When you do that you get that angle A is 183 - 4x. Angle B, subtracted from 180 is 168 - 9x. So in order to find out what x is, you have to add all the angles together and set them equal to 180, which is the measure of all the angles of the triangles, right?

OpenStudy (imstuck):

That gives you an equation of 183 - 4x + 168 - 9x + 6x + 3 = 180. Combining like terms you can solve for x. 183 + 168 + 3 = 354. But you have to move it over to the other side of the equal sign and 180 - 354 = -174. But it's ok that it is negative, because when you combine your x's you get -7x. -7x = -174. x = 24.8 or if you round up, 25.

OpenStudy (anonymous):

Ohhh I see! So we need to change x into 25 to get their angles

OpenStudy (imstuck):

YES!!!! You go!

OpenStudy (anonymous):

6x+3=153???

OpenStudy (anonymous):

6(25)+3=153?

OpenStudy (imstuck):

Oops, except its not coming out right. Let me look at it some more.

OpenStudy (imstuck):

I think figured out my error...give me one more sec...I didn't go anywhere.

OpenStudy (imstuck):

OMG I am totally stumped on this one. I keep getting x = -3 now, and that's not possible. I am going to keep working on it. I'll let you know if I come up with anything lucrative.

OpenStudy (imstuck):

I keep getting negative angles all over the place with this one; don't know what to tell you except that angles cannot have negative measures.

OpenStudy (imstuck):

Is the 6x + 3 the measure of an angle or the length of a side?

OpenStudy (anonymous):

Sorry for afk. They're all angles

OpenStudy (anonymous):

We can imagine a parallel line above and then relate the angles and using algebra and that the angles sum on each corresponding line is equal to 180 we can hopefully solve for X